Please help me simplify this equation: (7d^3/2 * 2g^5/6)(2g^3/2 * 7d^5/6)
What I have so far is ( 7*2*2*7)=196. I'm confused with the rest. Any help will be so appreciated! Thank you!

Answer by rothauserc(4718)

Add exponents of like bases
:
d^(3/2) * d^(5/6) = d^(14/6) = d^(7/3)
g^(5/6) * g^(3/2) = g^(14/6) = g^(7/3)
:
196*d^(7/3)*g^(7/3)
: